which is the correct way to Simpify 7(5-2m)-m
35-15m or 35-13m
thank you

Found 2 solutions by stanbon, jenni.fowler021510:
Answer by stanbon(50447)

Simpify 7(5-2m)-m
35-15m or 35-13m
---------------------------
7(5-2m)-m
---
= 35-14m-m
----
= 35-15m
=================
Cheers,
Stan H.

Answer by jenni.fowler021510(3)

The correct answer is 35-15m
reason: 7(5-2m)-m
7*5=35
7(-2)=-14
-14m-1m=-15m
and them put back together
35-15m